Solve the unknown value




Force Mass Acceleration 1. 45 kg 7.6 m/s2

2. 255 N 11.25 m/s2

3. 546 N 35 kg

4. 45 N 11 m/s2

5. 85 kg 8 m/s2

1
Expert's answer
2021-09-03T12:12:16-0400

1

"F=ma=1.45\:\rm kg*7.6\: m/s^2=11\: N"

2

"m=\frac{F}{a}=\frac{255\:\rm N}{11.25\:\rm m/s^2}=22.7\:\rm kg"

3

"a=\frac{F}{m}=\frac{546\:\rm N}{35\:\rm kg}=16\:\rm m/s^2"

4

"m=\frac{F}{a}=\frac{45\:\rm N}{115\:\rm m/s^2}=4.1\:\rm kg"

5

"F=ma=85\:\rm kg*8\: m/s^2=700\: N"